A body of mass m is accelerated uniformly from rest to a speed v in a time t.The instantaneous power deliverd to body as a function of time , is given by

Grade:11

2 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
We can find the acceleration produced by the mass as given by:
a=\frac{v}{t}
Hence, the instantaneous value of power would be given by:
P=\frac{Mv}{t}\times v=\frac{Mv^{2}}{t}
Regards
Sumit
Aman Sharma
6 years ago
Instantaneous power= force*displacement/time^2So force= m.aTherefore, power= m.a.s/t^2 a.s= v^2/2Hence power = 1/2 mv^2/t^2